# Learning-based Multi-View Stereo: A Survey

### v2

@@ 6 Discussions
In this section, we summarize and discuss the performance of learning-based MVS methods, including supervised online methods with depth estimation (Sec. 6.1), supervised offline methods with depth estimation (Sec. 6.2), unsupervised methods with depth estimation (Sec. 6.3) and methods without depth estimation (Sec. 6.4). Moreover, we discuss potential directions for future research (Sec. 6.5).
TABLE I: Quantitative results of supervised online MVS methods on ScanNet [26] and 7-Scenes [179].
@@
abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow
@@
TABLE II: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for unsupervised depth map-based MVS methods.“Finetuned with BlendedMVS” denotes whether the methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22].
Methods DTU [21] Finetuned with Tanks and Temples [22]
Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [24] Intermediate $F_{1}$uparrow Advanced $F_{1}$uparrow
End-to-end Unsup_MVSNet [127] 0.881 1.073 0.977 - -
MVS2 [125] 0.760 0.515 0.637 37.21 -
M3VSNet [126] 0.636 0.531 0.583 37.67 -
JDACS-MS [41] 0.398 0.318 0.358 45.48 -
RC-MVSNet[42] 0.396 0.295 0.345 55.08 30.82
ElasticMVS[131] 0.374 0.325 0.349 57.88 37.81
CL-MVSNet[132] 0.375 0.283 0.329 59.39 37.03
Multi-stage Self-supervised CVP-MVSNet [128] 0.308 0.418 0.363 43.48 -
U-MVS [43] 0.354 0.354 0.354 57.15 30.97
KD-MVS [44] 0.359 0.295 0.327 64.14 37.96
TABLE III: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for supervised offline MVS methods. “Finetuned with BlendedMVS” denotes whether methods are finetuned on BlendedMVS dataset [24] before evaluation on Tanks and Temples [22] and ETH3D [23].
@@
Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [24] Intermediate $F_{1}$uparrow Advanced $F_{1}$uparrow Training $F_{1}$uparrow Test $F_{1}$uparrow
@@
6.1 Supervised Online MVS with Depth Estimation
Typically, online MVS methods are trained on ScanNet [26] and then evaluated on ScanNet [26] and 7-Scenes [179]. We summarize the quantitative results of online MVS methods in Tab. I. By incorporating temporal information in depth estimation,  [60, 188, 30] achieve better performance than MVDepthNet [25]. However, this usually increase the network complexity. Instead of temporal fusion, SimpleRecon [34] injects cheaply available metadata into the cost volume and achieves state-of-the-art performance on both ScanNet and 7-Scenes.
6.2 Supervised Offline MVS with Depth Estimation
Offline MVS methods are usually trained on DTU [21] and then evaluated on DTU [21], Tanks and Temples [22] and ETH3D [23]. Recently, many methods further finetune the DTU-pretrained model on BlendedMVS [24] before evaluating on Tanks and Temples and ETH3D since it contains large-scale scenes.
6.2.1 Benchmark Performance
Tab. III summarizes quantitative results of offline MVS methods. Compared with those methods that use direct 3D CNN for regularization, the methods that use RNN and coarse-to-fine regularization perform much better. Since RNN based methods have bad time efficiency, we can find that coarse-to-fine methods become main stream of the community because of their impressive performance and high efficiency in both memory and run-time. In addition, iterative methods based on traditional PatchMatch [39, 102] or RAFT [40, 104, 103] also become popular since they achieve comparable performance as state-of-the-art methods with more lightweight structure. For benchmarks, most learning-based methods perform explicitly better than traditional methods on DTU and Tanks and Temples. However, on ETH3D, traditional methods [189, 123] perform better than all existing learning-based methods. We conjecture that the strong viewpoint variations and large textureless regions in human-made environments still make learning-based methods struggle. Therefore, it is valuable to pay more attention to ETH3D to evaluate the robustness of learning-based methods in real-world scenes.
6.2.2 Memory Consumption and Run-time
Low run-time and memory consumption are crucial in most industrial applications with limited computational power and storage, e.g., robotics and AR/VR. Because of the low-resolution input and the simplicity of network structures, online MVS methods [25, 30, 34] usually achieve high efficiency in both memory and run-time. In contrast, offline MVS methods focus on high-resolution images and usually use computationally expensive network modules, e.g., 3D convolutions, to improve reconstruction quality. These explicitly increase run-time and GPU memory. Recently, many researchers tried to improve the efficiency while maintaining the reconstruction accuracy as other offline MVS methods. We compare the efficiency of state-of-the-art offline MVS methods [31, 39, 104, 40, 100, 79, 111, 76, 86] in Fig. 5. All these methods share similar coarse-to-fine structures to improve efficiency. However, we observe that some methods [31, 79, 111, 76] still require lots of computation resource and are not efficient in either memory or run-time. To further reduce run-time and memory, [39, 104, 40, 100] carefully simplify the architecture, e.g., replacing costly 3D convolution with 2D convolution. Moreover, these methods still achieve competitive performance when compared with the top-performing methods. Therefore, it is promising to further improve both the performance and efficiency for practical applications.
Figure 5: GPU memory consumption and run-time of state-of-the-art offline methods [31, 39, 104, 40, 100, 79, 111, 76, 86] on DTU dataset [21]. The image resolution and the number of images are set to $1600\times 1152$ and 5 respectively. For fair comparison, experiments are done on one workstation with a NVIDIA 2080 Ti GPU.
6.3 Unsupervised MVS with Depth Estimation
We summarize the results of unsupervised MVS methods in Tab. II. Based on the conventional photometric consistency loss (Sec. 4.1), unsupervised MVS methods may have difficulties with many challenging situations, e.g., varying lighting conditions or occlusions, and suffer from downgraded reconstruction results. Therefore, many methods propose to employ different strategies to improve the robustness, e.g., segmentation consistency [41], optical flow consistency [43], neural rendering consistency [42], pseudo-label generation with geometric filtering [128, 44], multi-iteration training [128, 44] and featuremetric loss [44]. It is inspiring that KD-MVS [44] already outperforms its backbone, CasMVSNet [31], and many other fully supervised methods. However, since unsupervised methods mainly test on relatively simple scenes [21, 22], their scalability to more complex and large-scale scenes, e.g., ETH3D [23], remains a question. This good generalization capability across various scenes and datasets is crucial for practical applications. Therefore, evaluating unsupervised MVS methods on various scenes is an important task for future research. Most unsupervised methods use simple MVS backbones [15, 31, 75] and mainly focus on the training strategy. Employing state-of-the-art architectures is also a potential direction that can further boost the performance.
6.4 Learning-based MVS without Depth Estimation
6.4.1 Voxel-based Methods
We summarize the quantitative results on ScanNet [26] in Tab. IV. Since the voxel representation increases the computation overhead explicitly when the scale of scene increases, voxel-based methods mainly focus on indoor scenes, e.g., ScanNet [26]. We find that depth map-based MVS methods, e.g., SimpleRecon [34], fuse depth maps with traditional TSDF fusion [65] and can achieve comparable performance as these voxel-based methods with lower complexity. This further shows the advantages of depth map-based methods over voxel-based methods.
TABLE IV: Quantitative results of voxel-based MVS methods on ScanNet [26].
Methods Prec. $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
Atlas [46] 0.675 0.605 0.636
NeuralRecon [47] 0.630 0.612 0.619
TransformerFusion [134] 0.728 0.600 0.655
VoRTX [135] 0.767 0.651 0.703

### v3

@@ 6 Discussions
In this section, we summarize and discuss the performance and efficiency of learning-based MVS methods. Moreover, we discuss potential directions for future research.
TABLE II: Quantitative results of point cloud evaluation on Tanks & Temples [23]. F-score is listed (the higher, the better). MapAnything [63] reconstructs point clouds directly from unposed images. All images are used for MapAnything, except for Courthouse (50% subset due to memory limitations).
Representation Methods Barn Caterpillar Courthouse Ignatius Meetingroom Truck Mean
Depth map DeepVideoMVS [31] - - - 0.01 - 0.01 -
CasMVSNet [32] 0.52 0.42 0.36 0.71 0.35 0.59 0.49
$D^{2}$HC-RMVSNet [37] 0.64 0.57 0.42 0.74 0.33 0.70 0.57
AA-RMVSNet [89] 0.61 0.58 0.40 0.77 0.36 0.71 0.57
IterMVS [40] 0.62 0.49 0.38 0.78 0.38 0.68 0.55
TransMVSNet [91] 0.57 0.47 0.38 0.73 0.42 0.63 0.53
MVSFormer++ [93] 0.66 0.59 0.42 0.78 0.40 0.69 0.59
CasDiffMVS [119] 0.69 0.54 0.42 0.81 0.45 0.70 0.60
Voxel NeuralRecon [47] - 0.01 - 0.07 - - -
VisFusion [192] - 0.09 - 0.09 0.03 0.15 -
NeRF NeuS [49] 0.29 0.29 0.17 0.83 0.24 0.45 0.38
Geo-NeuS [160] 0.33 0.26 0.12 0.72 0.20 0.45 0.35
Neuralangelo [50] 0.70 0.36 0.28 0.89 0.32 0.48 0.50
3DGS SuGaR [55] 0.14 0.16 0.08 0.33 0.15 0.26 0.19
2DGS [56] 0.36 0.23 0.13 0.44 0.16 0.26 0.30
GOF [193] 0.51 0.41 0.28 0.68 0.28 0.58 0.46
PGSR [57] 0.66 0.41 0.21 0.80 0.29 0.60 0.50
Feed-forward Point-based MapAnything [63] 0.14 0.34 0.10 0.18 0.12 0.23 0.19
6.1 Performance Comparison
To provide a thorough comparison across different methods, we evaluate state-of-the-art methods on Tanks & Temples [23] and summarize the results in Tab. II.
For depth map-based methods, we find that online methods [31] perform worse than offline methods [32, 91, 119, 93]. Online methods typically sample sparse depth hypotheses from a pre-defined depth range to construct a lightweight cost volume, which limits the accuracy in large-scale scenes. In addition, online methods are limited by causal/temporal processing and cannot exploit geometric constraints from future frames as offline methods.
For voxel-based methods, we observe that they fail on both outdoor and indoor scenes. From the representation’s perspective, we can attribute the performance gap to the following two aspects. (a) Voxel-based methods usually specify the spatial location of the potential voxels given the camera trajectory, with a metric voxel resolution. The inherent memory complexity and limited resolution of voxel grids are fundamental bottlenecks for large-scale, unbounded scenes. (b) Unlike depth-based representations, which estimate geometry primarily based on photometric consistency on the 2D image plane and allow for sharp discontinuities (suitable for complex occlusions in outdoor scenes), voxel-based SDFs impose a global 3D continuity constraint. While this constraint helps regularize geometry in texture-less indoor regions, it acts as a limiting factor in large-scale scenes by enforcing smoothness where sharp edges or fine details are expected.
For NeRF/3DGS-based methods, we observe that they achieve high accuracy in small scenes, e.g., Ignatius, producing high-quality reconstruction with fine-grained details. However, the performance is notably worse than offline methods on large scenes, e.g., Courthouse. Scaling NeRF/3DGS to large-scale scenes is still an open topic in research.
For large feed-forward point-based methods, although the reconstructed point clouds appear reasonable, their quantitative accuracy is typically low. This is due to several challenges: (1) they must estimate camera poses, significantly increasing task complexity; however, models like DUSt3R skip explicit camera projection, and VGGT predicts camera tokens in parallel, bypassing camera geometry during the forward pass. This flexibility in handling unposed data, while advantageous, means that metric evaluation is highly sensitive to pose alignment and can limit accuracy. (2) These models are typically not trained on scenes with hundreds of views, creating an out-of-distribution problem for larger datasets; and (3) huge memory consumption compels the use of low-resolution images or fewer input frames, directly limiting the fidelity of the final reconstruction.
6.2 Efficiency Comparison
Low run-time and memory consumption are crucial in most industrial applications with limited computational power and storage, e.g., robotics and AR/VR.
Compared to other representations, depth-map based MVS is usually more suitable for these applications. Volume-based methods are difficult to scale to large scenes due to substantial memory storage. NeRF-based and 3D Gaussian Splatting-based methods typically require per-scene optimization for high-quality reconstruction. Large feed-forward methods have huge backbones and thus their GPU memory consumption is high. We compare the efficiency of state-of-the-art depth-map based MVS methods [32, 39, 31, 112, 40, 108, 91, 120, 88, 194] in Fig. 8. Although online methods [31] achieve high efficiency in run-time with low-resolution images, we observe that their run-time and GPU memory consumption are higher than many offline methods [39, 112, 40, 108, 119] with high-resolution images. To improve efficiency, offline methods [39, 112, 40, 108, 119] carefully simplify the architecture, e.g., replacing costly 3D convolution with 2D convolution. Moreover, these methods, especially CasDiffMVS [119], achieve competitive performance when compared to the top-performing methods [194, 93]. Therefore, it is promising to further improve both the performance and efficiency for practical applications.
Figure 8: GPU memory consumption, run-time, and reconstruction error of state-of-the-art depth map-based methods [31, 39, 112, 40, 108, 91, 120, 194] on DTU [22]. To fairly compare efficiency, we fix the image resolution and the number of input images, which are set to $1600\times 1152$ and 5 respectively. All experiments are done on the same workstation with a NVIDIA 2080 Ti GPU.
6.3 Future Research Directions
6.3.1 Datasets and Benchmarks
For MVS methods, ScanNet [27] and DTU [22] are two main training datasets, while ScanNet [27], DTU [22], Tanks & Temples [23] and ETH3D [24] are the main evaluation benchmarks.
For training, the scene scale of ScanNet and DTU is relatively small (room-scale for ScanNet and object-scale for DTU) and their quality is not satisfactory [195]. Recently, many researchers have tried to improve the dataset quality. For example, TartanAir dataset [196] is a large-scale synthetic dataset that is collected in photo-realistic simulation environments. BlendedMVS [25] introduces more large-scale scenes and improves the performance of many MVS methods on real-world scenes. MVImgNet [197] and Objaverse [198] are large datasets of multi-view images collected by shooting videos of real-world objects or rendering 3D object models. However, both of them focus on object-level scenes, whereas large-scale scenes are essential for many applications. DL3DV [199] attempts to address this issue by capturing videos of large-scale real-world scenes. However, it lacks the ground truth geometry required for MVS training. Therefore, improving the scalability and quality of datasets is an important research direction. One direction is to use advanced sensors, e.g., depth camera and LiDAR, to capture high-quality data in diverse real-world scenario [78, 200]. The other direction is to exploit large-scale video data, e.g., Youtube videos, and extract high-quality pseudo ground-truth with state-of-the-art models.
For evaluation, the main benchmarks [27, 22, 23, 24] are all introduced before 2018. Some of these benchmarks are currently saturated and the methods are difficult to further improve the performance on them. Therefore, it is meaningful to introduce new benchmarks to evaluate the robustness and performance of learning-based methods in challenging real-world scenes. For example, the new benchmarks are expected to include challenging settings, such as dynamics, low-texture, reflections, and appearance changes [201].
6.3.2 View Selection
As discussed in Sec. 2.2, view selection is crucial for triangulation quality. Picking neighboring views that are suitable for triangulation can not only improve the reconstruction accuracy but also reduce useless computation for the bad views, e.g., views with strong occlusions. However, view selection is often overlooked and not well studied. For example, all offline depth map-based MVS methods [32, 39, 91, 108] follow MVSNet [9] and use the same simple heuristic strategy to compute scores for neighboring views and sort them. Online depth map-based MVS methods [67, 31, 34] and voxel-based methods [47] adopt heuristic strategies to choose views with enough pose-distance. Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design. Current view selection strategies simply pick the same set of neighboring views for all the pixels in the reference image, and it is probable that different reference pixels have different optimal choices of neighboring views. The learned pixel-wise view weight [101, 39, 40] addresses this issue by assigning different weights to source views for each pixel based on their visibility across neighboring views. For large feed-forward methods such as VGGT [62], cross-view attention provides an alternative mechanism for adaptive view selection with the attention weights.
6.3.3 Feature Extraction
Currently, most of the depth map-based [9, 31] and voxel-based [47] methods mainly use 2D CNN and FPN [83] structure to learn multi-scale features. To flexibly improve the receptive fields, Deformation Convolution [202] and attention mechanism [144] are applied  [98, 90, 89, 39, 91]. However, they still cannot generalize well in challenging regions, such as reflective and low-textured areas. Motivated by the success of vision transformer [203], recent works [204, 93, 61, 62] introduce pre-trained vision transformers, e.g., ViT [203] and DINO [94, 187, 205], to capture global information, alleviating the issues of reflections and texture-less regions. However, vision transformers are usually trained and tested on low-resolution images due to their complexity. To accurately reconstruct the scene with high-resolution inputs, it is important to utilize fine-grained information for precise feature matching across views, which can be nicely captured with CNN. Therefore, an interesting direction is to combine vision transformer and CNN in a coarse-to-fine framework, where vision transformer is used in low-resolution and CNN is used in high-resolution for local refinement.
6.3.4 Generative Reconstruction
MVS tends to be sensitive to occlusions and sparse viewpoints because of their strict geometric requirements. However, these challenging conditions are quite common in practice, such as autonomous driving. To this end, while mainstream MVS methods follow a regression-based paradigm, it would be promising to leverage the inpainting power of generative models to produce reasonable reconstruction in unseen regions. For example, ReconFusion [206] explores to distill the generated novel views to a NeRF representation for sparse-view reconstruction. Difix3D+ [207] refines noisy 2D renderings from Gaussian Splatting with reference views and Stable Diffusion [208]. Achieving visually pleasing 3D reconstruction while maintaining the reconstruction fidelity and geometric consistency with generative models is a significant research trend, which will also power the development of world models.
6.3.5 Efficiency
Though reconstruction accuracy is considered the most important factor in MVS, it is also critical to have models that can run in near real-time and with low memory, e.g., AR/VR and autonomous driving. Comparatively, depth map-based methods are most suitable for achieving efficiency because of their conciseness, flexibility, and scalability. Though online MVS methods can achieve real-time estimation with images of low resolutions, they may have issues with memory consumption since many large backbones are usually used for feature extraction [31, 34]. In addition, the efficiency in both run-time and memory will drop when image resolution increases, as shown in Fig. 8. For offline MVS methods, as discussed in Sec. 6.2, some recent methods try to improve efficiency with carefully simplified network architectures. However, this usually limits the performance when compared to state-of-the-art methods. In addition to designing lightweight modules [119], using model compression techniques and knowledge distillation are potential directions to not only keep the good performance but also improve efficiency.
Large feed-forward models heavily rely on attention mechanism to capture intra-frame and inter-frame relations. However, the high computational cost restricts the use of high-resolution or numerous input images, which can degrade performance [62, 63]. To reduce computational cost of attention mechanism, one direction is to introduce token merging to reduce attention redundancy [209].
6.3.6 Prior Assistance
MVS mainly relies on assessing local photometric consistency to find the optimal matching across images. Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made scenes [24]. Therefore, using prior information to guide the MVS algorithm in these challenging regions is a promising research direction. We elaborate 4 typical examples of prior assistance as follows.
Monocular Depth: Recently, significant progress has been made in monocular depth estimation [210, 181], with existing methods achieving accurate results on diverse in-the-wild datasets. Leveraging monocular depth as a strong prior can enhance robustness in challenging scenarios such as low-textured regions and reflective surfaces [161, 59]. For example, monocular depth can be used to regularize or inpaint depth estimation in these challenging regions.
Surface Normal: As a non-local representation of the geometry compared with the depth map, surface normal has been proven effective in improving reconstruction performance in recent works [211, 212]. Similar to monocular depth [210, 181], monocular normal estimation [163] produces high-quality normal priors on in-the-wild data, which can improve reconstruction quality in low-textured regions[161, 213].
7 Conclusion
Supplementary Material
1 Preliminaries
1.1 Plane Sweep
We take a toy example of three cameras viewing the same object as illustrated in Fig. 9. Plane sweep is then performed on the frustum of cam_x (termed as the reference camera) by creating a series of fronto-parallel hypothesized planes within a given range (typically $d_{\textrm{min}},d_{\textrm{max}}]$), with each plane corresponding to a depth value w.r.t. cam_x. Let’s then examine two 3D points, $M$ and $M^{\prime}$, as examples of occupied and unoccupied positions on hypothesized planes, respectively. For $M$, all of the three cameras capture the identical point lying on the geometry surface with photometric consistency. In contrast, the photometric consistency of the observations of $M^{\prime}$ are poor, indicating that $M^{\prime}$ is an invalid hypothesis. To evaluate the similarity of the observations towards $M$ and $M^{\prime}$, the images of cam_1 and cam_2 (termed as the source cameras) are warped to each plane by homography. These warped images are then compared against the scaled images of cam_x. Depth hypotheses with high similarity measures are considered reliable.
Figure 9: Illustration of plane sweep algorithm [33]. To estimate the depth map for reference image (cam_x), neighboring source images (cam_1, cam_2) are projected with homography to fronto-parallel planes of the reference view frustum.
Regarding selecting $D$ depth hypotheses from the triangulated depth range $d_{\textrm{min}},d_{\textrm{max}}]$, there are two main schemes, namely the forward depth sampling and the inverse depth sampling. The naive forward sampling divides the depth range into D1D-1 depth intervals with identical lengths. Given a depth index $k$, we have
$d_{k}=d_{\textrm{min}}+\frac{k}{D-1}(d_{\textrm{max}}-d_{\textrm{min}}),k=0,\ldots,D-1,$ (25)
where the depth hypotheses distribute uniformly between the two ends. While for the inverse sampling scheme [35], we sample uniformly in the multiplicative inverse of $d$, such that
$\frac{1}{d_{k}}=\frac{1}{d_{\textrm{max}}}+\frac{k}{D-1}(\frac{1}{d_{\textrm{min}}}-\frac{1}{d_{\textrm{max}}}),k=0,\ldots,D-1.$ (26)
In this way, the distribution of sampled depth values becomes sparser with $d$ approaching to $d_{\textrm{max}}$. When reconstructing unbounded outdoor scenes, where the depth range is rather large, the inverse sampling will be a reasonable choice since it samples more densely at the foreground. In addition, as revealed in [101], the inverse sampling leads to a uniform sampling on the projected epipolar lines of source images.
1.2 Depth Fusion
To remove outliers and improve reconstruction quality, offline MVS methods filter out outliers before converting the depth maps to point clouds. Depth filtering mainly involves photometric and geometric filtering [9].
For photometric consistency filtering, a per-pixel confidence is estimated to measure the confidence of depth estimation, i.e., probability that the ground truth depth is within a small range near the estimation. Then a threshold can be set to filter depth values with low confidence. To estimate this confidence, existing methods either follow MVSNet [9] and estimate photometric confidence from the probability volume [36, 37, 89, 9, 32, 39] or use a network to predict the confidence of estimation [40, 119].
For geometric consistency filtering, the consistency of depth estimations are measured among multiple views. For a pixel $\mathbf{p}$ in the reference view 0, we project it to pixel $\mathbf{p}^{\prime}$ in its $i$-th neighboring view through its depth prediction $d_{0}(\mathbf{p})$. After looking up the depth for $\mathbf{p}^{\prime}$, $d_{i}(\mathbf{p}^{\prime})$, we re-project $\mathbf{p}^{\prime}$ back to the reference view at pixel $\mathbf{p}^{\prime\prime}$ and look up its depth $D_{0}(\mathbf{p}^{\prime\prime})$. We consider pixel $\mathbf{p}$ and its depth as consistent to the $i$-th neighboring view, if the distances, in image space and depth, between the original estimate and its re-projection satisfy:
$\xi_{d}=\|\mathbf{p}-\mathbf{p}^{\prime\prime}\|_{2}\leq\tau_{1},$ (27)
$\xi_{p}=\frac{\|d_{0}(\mathbf{p}^{\prime\prime})-d_{0}(\mathbf{p})\|}{d_{0}(\mathbf{p})}\leq\tau_{2},$ (28)
where $\tau_{1}$ and $\tau_{2}$ are thresholds. The pixels are considered to be reliable estimations if they are consistent in at least $N$ neighboring views. Recently, instead of using predefined thresholds for $\tau_{1}$ and $\tau_{2}$, dynamic consistency checking [37] is proposed to dynamically aggregate geometric matching error among all views and improve the robustness of reconstruction. Specifically, the dynamic multi-view geometric consistency $C_{\textrm{geo}}(\mathbf{p})$ is computed as:
$C_{\textrm{geo}}(\mathbf{p})=\sum_{i}\exp{(-(\xi_{p}+\lambda\xi_{d}))},$ (29)
where $\lambda$ is a weight to balance the reprojection error in two metrics. Then the outliers with $C_{\textrm{geo}}(\mathbf{p})$ smaller than a threshold are filtered.
1.3 Evaluation Metrics
1.3.1 2D Metrics
We list the commonly used metrics as follows: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel) and inlier ratio with threshold 1.25 ($\delta<1.25$):
$\text{Abs}=$ $\frac{1}{N}\sum_{\mathbf{p}}|d(\mathbf{p})-\hat{d}(\mathbf{p})|,$ (30)
$\text{Abs Rel}=$ $\frac{1}{N}\sum_{\mathbf{p}}\frac{|d(\mathbf{p})-\hat{d}(\mathbf{p})|}{d(\mathbf{p})},$
$\text{Inlier Ratio}=$ $\frac{1}{N}\sum_{\mathbf{p}}\mathbbm{1}\left[\frac{d(\mathbf{p})}{1.25}<\hat{d}(\mathbf{p})<1.25d(\mathbf{p})\right],$
where $d(\mathbf{p})$ denotes the ground truth depth, $\hat{d}(\mathbf{p})$ denotes the estimation, $N$ denotes the number of pixels with valid depth measurements and $\mathbbm{1}[\cdot]$ denotes the indicator function.
1.3.2 3D Metrics
Precision/Accuracy: Precision/Accuracy measures the percentage of predicted points that can be matched to the ground truth point cloud. Considering a point $\mathbf{P}_{p}$ in the predicted point cloud, it is considered to have a good match in the ground truth point cloud $\{\mathbf{P}_{g}\}$ if
$\|\mathbf{P}_{p}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{g}\}}\|\mathbf{P}-\mathbf{P}_{p}\|_{2}\|_{2}\leq\tau,$ (31)
where $\tau$ is a scene-dependent threshold assigned by datasets, usually set to a large value for large-scale scenes. Note that in some datasets, instead of being measured by percentage [23, 24], precision/accuracy is measured by mean or median absolute distance [22].
Recall/Completeness: Recall/Completeness measures the percentage of ground truth points that can be matched to the predicted point cloud. For a point $\mathbf{P}_{g}$ in the ground truth point cloud, it is considered a good match in the predicted point cloud $\{\mathbf{P}_{p}\}$ if
$\|\mathbf{P}_{g}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{p}\}}\|\mathbf{P}-\mathbf{P}_{g}\|_{2}\|_{2}\leq\tau.$ (32)
Similar to precision/accuracy, recall/completeness can be measured by the percentage of points [23, 24] or measured by the absolute distance [22], similar to Chamfer distance.
F-Score: The two aforementioned metrics measure the accuracy and completeness of predicted point clouds. However, each of these metrics alone cannot present the overall performance since different MVS methods adopt different assumptions. A stronger assumption usually leads to higher accuracy but lower completeness. If only precision/accuracy is reported, it would favor MVS algorithms that only include estimated points of high certainty. On the other hand, if only recall/completeness is reported it would favor MVS algorithms that include everything, regardless of accuracy. Therefore, F-score, an integrated metric, is introduced [23, 24]. F-score is the harmonic mean of precision and recall. It is sensitive to extremely small values and tends to get more affected by smaller values, which means that F-score does not encourage imbalanced results. However, in most cases, F-score still suffers from unfairness due to the limitations of ground truth, e.g., sparse and incomplete point clouds may penalize for filling in the areas that are not present in the ground truth [46].
2 Supervised Method with Depth Representation
2.1 Cost Volume Regularization
In Fig. 10, we visualize the three main categories to perform cost volume regularization for offline MVS methods: direct 3D CNN, coarse-to-fine and RNN.
Figure 10: Illustration of typical cost regularization schemes for offline MVS methods that use 4D cost volumes. (a) direct 3D CNN regularization [9, 104, 105, 35] applies a 3D CNN to aggregate contextual information; (b) coarse-to-fine regularization [32, 38, 87, 101] constructs multi-scale cost volumes based on coarse prediction and uses 3D CNN regularization on each scale; (c) RNN regularization [36, 37, 89] sequentially regularizes 2D slices of the cost volume to reduce memory consumption.
TABLE III: Quantitative results of voxel-based MVS methods on ScanNet [27].
Methods Prec. $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
Atlas [46] 0.565 0.598 0.578
NeuralRecon [47] 0.684 0.479 0.562
TransformerFusion [48] 0.648 0.547 0.591
VoRTX [142] 0.688 0.607 0.644
VisFusion [192] 0.695 0.527 0.598
3 Benchmark Performance
In this section, we summarize the performance of different methods on the common benchmarks they use. Specifically, we summarize the quantitative results of online depth map-based MVS in Tab. IV, offline depth map-based MVS in Tab. V, voxel-based MVS in Tab. III, and NeRF/3D Gaussian Splatting-based MVS in Tab. VI.
TABLE IV: Quantitative results of online MVS methods on ScanNet [27] and 7-Scenes [76].
@@ 3 Benchmark Performance
abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$ $\uparrow$ abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$ $\uparrow$
@@
TABLE V: Quantitative results of point cloud evaluation on MVS benchmarks [22, 23, 24] for offline MVS methods. “Finetuned with BlendedMVS” denotes whether methods are finetuned on BlendedMVS dataset [25] before evaluation on Tanks and Temples [23] and ETH3D [24].
@@
Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [25] Intermediate $F_{1}$ $\uparrow$ Advanced $F_{1}$ $\uparrow$ Training $F_{1}$ $\uparrow$ Test $F_{1}$ $\uparrow$
@@
DiffMVS [119] 0.318 0.297 0.308 63.39 39.69 74.86 82.10
CasDiffMVS [119] 0.310 0.286 0.298 65.87 41.81 76.76 85.11